OpenStudy (anonymous):

How many solutions are there to the equation below? 3x - 10(x + 2) = 13 - 7x

OpenStudy (anonymous):

3x -10x -20 = 13 -7x -20 = 13 no solutions

OpenStudy (anonymous):

that is inccorect Coolsector.